OpenStudy (anonymous):

Find the vertex of the quadratic equation: y = 2x2 + 8x − 4.

OpenStudy (anonymous):

You need to complete the square. Do you know how to do it?

OpenStudy (anonymous):

the vertex is (2,-12)

OpenStudy (anonymous):

(-2, -12)

OpenStudy (anonymous):

can you explain how yuo got that?

OpenStudy (anonymous):

not sure how to do this

OpenStudy (anonymous):

2(x^2 + 4x - 2) = y x^2 + 4x - 2 = y/2 x^2 + 4x +4 = y/2 + 6 (x+2)^2 = y/2 + 6 (x+2)^2 - 6 = y/2 2(x + 2)^2 - 12 = y so the vertex is (-2, -12)

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

you can use the equation -(b/2a) to find the vortex 2x^(2) + 8x − 4. where a=2 b = 8 so -(8/2(2)) = -2 = x now sub -2 into the equation to find the y value and you have your vortex